# On Selection of Cross-Section Averages in Non-stationary Environments

### v3

@@ 2 Econometric Setup
$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\tau b\ln(T)+\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}})\right)+R_{N,T},$ (2.5)
@@
where $|R_{N,T}|=o_{p}(1)$ is the remainder, $\lambda_{j}(\mathbf{A})$ is the $j$-th eigenvalue of $\mathbf{A}$, while $b$ is a number of strictly positive eigenvalues. Because $b$ is finite ($b\leq k$), it is clear that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to+\infty$ at a rate of $O_{p}(\ln(T))$. Two comments are in order. Firstly, in the Supplement we show that $|R_{N,T}|=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$, and it is not necessarily positive. This means that for $\tau\approx 1$ and a small $N$, the remainder vanishes slowly as $T\to\infty$. Hence, it may be necessary to rely on large $N,T$ combinations to detect the minimum of IC outside of $M\subset M_{0}$ region if $\mathbf{F}$ is very persistent. We explore (and confirm) such risks in Monte Carlo simulations in Section 3. Secondly, while this result is similar to Lemma A3 in Bai, (2004), there the divergence rate is $O_{p}(T/\ln\ln(T))$, since the law of iterated logarithm is used to arrive at the expression analogous to (2.5) when factors have unit root. Theorem 1 leads to our consistency result.
Theorem 1. Under conditions of Proposition 1 with $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$, we have
$\mathbb{P}\left(\widehat{M}=M_{0}\right)\to 1$
as $(N,T)\to\infty$.
@@
Proof. Follows directly from Proposition 1, because $\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})>0\right)\to 1$ under (a)-(b) and the penalty properties.
@@
The key message of Theorem 1 is that the IC of DVS (and MW) is consistent, but exhibits hybrid asymptotic properties characteristic to stationary and non-stationary factors. Consequently, two forces are in effect: $p_{N,T}$ should still be negligible, but the factor estimation error can be “large” if $M\subset M_{0}$ as indicated in (2.5). To balance these properties in practice, it may be tempting to further experiment with the penalty. In an analogous theorem of Bai, (2004), it is required that $p_{N,T}\to\infty$ and $p_{N,T}/\ln(T)\to 0$ if the criterion is based on a logarithmic-valued sum of squared residuals as is the case here (see footnote 3 in that study). Importantly, this condition does not interfere with consistency in our Theorem 1, because clearly $p_{N,T}=o(\ln(T))$ still holds. Therefore, we compare both types of penalties in the simulations to illustrate the outcome if practitioners simply extrapolate recommendations from the PC literature. We let $\widetilde{p}_{N,T}=\ln(T)p_{N,T}$, which follows the suggestion in (12) in Bai, (2004), and obeys $\widetilde{p}_{N,T}\to\infty$ and $\widetilde{p}_{N,T}/\ln(T)\to 0$.
@@
Remark 1. Note that under stationarity, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}c>0$ when $M\subset M_{0}$, according to the results in DVS. Our Proposition 1 naturally accommodates this result under $\tau=0$, because, by following the same approximation steps, (2.5) becomes
@@
$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\sum_{j:\lambda>0}\ln\left(1+\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}})\right)+R_{N,T}\to_{p}c>0$
@@
as $(N,T)\to\infty$ because each summand is greater than 1 w.p. 1. Also, $|R_{N,T}|=O_{p}(N^{-1})+O_{p}((NT)^{-1/2})$, which is the same as in DVS under $\tau=0$, as expected from the discussion of (2.5).
@@ 3 Simulation Study
where $\mathbf{X}_{i}=\left[\mathbf{x}_{i,1},...,\mathbf{x}_{i,T}\right]^{\prime}$ is a $T\times k$ matrix of observable regressors, $\boldsymbol{\beta}=\iota^{\prime}0.5$ a $k\times 1$ vector of unobserved parameters. The factor loadings are generated as $\boldsymbol{\Gamma}_{i}=[\mathbf{I}_{m},\mathbf{0}_{m\times(k-m)}]\psi_{i}$, that is, a matrix $m\times k$, with $\psi_{i}\sim N(1,1)$ and the elements of the vector $1\times m$ $\boldsymbol{\gamma}_{i}$ are drawn from $N(1,1)$. In the case of time- and space-independent errors, $\mathbf{e}_{i,t}$ is drawn from $N(\mathbf{0}_{k+1},\mathbf{I}_{k+1})$. In the case of errors correlated weakly across time and cross-section we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
@@
where $\boldsymbol{\epsilon}_{t}$ is a $N\times 1$ stack of $\epsilon_{i,t}\sim N(0,1)$ over $i$, and $\mathbf{w}_{i}$ is the $i$-th row ($1\times N$) of $\mathbf{W}$, a weight matrix with the $J$-th off diagonal elements equal to zero. $\rho$ is set to 0.5, $\kappa=0.2$ and $J=5$. The idiosyncratics in $\mathbf{X}_{i}$, $\mathbf{V}_{i}$, are also drawn as potentially weakly correlated over time and cross-sections. Here, $\mathbf{\Upsilon}_{t}$ is a $N\times k$ matrix, which stacks $\boldsymbol{\upsilon}_{i,t}\sim N(\mathbf{0}_{k},\mathbf{I}_{k})$ over $i$, and it is independent of $\boldsymbol{\epsilon}_{t}$. The vector $\boldsymbol{\iota}_{N,i}\in\mathbb{R}^{N\times 1}$ contains 1 in the $i$-th coordinate, and zeros elsewhere. $\rho_{v}=0$ for the time IID case and $\rho_{v}=0.5$ for weakly time-correlated idiosyncratics. In both cases, we allow for weak correlation across cross-sections, $\kappa_{v}=0.2$, $J_{v}=5$ and $\tilde{\mathbf{w}}_{i}$ is defined as $\mathbf{w}_{i}$ above. Therefore, we also test correlated idiosyncratics over time, whereas our theory prescribes uncorrelated innovations for $\tau\in(0,1)$. The factors are drawn as $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t}$, where
$\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=diag(g_{1},...,g_{m}),g_{j}\sim U(0,2)\quad\mathbf{u}_{f,t}\sim N\left(\mathbf{0}_{m},\sqrt{\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}}\right).$ (3.3)
@@
In the stationary factors case, the factors are drawn as a stable VAR with $\tau=0$. Two cases for non-stationary factors are considered: the non-stationary case of $\tau=0.4$ and the case close to local-to-unity process $\tau=0.9$. All simulations have 1000 repetitions and we compare DVO and MW. In the Supplement, we provide additional simulation evidence of the eigenvalue growth ratio (ER) based on Juodis, (2022) for comparison. The focus of this method is the number of factors ($m$), and it does not select the sufficient set of CAs. Nevertheless, its performance crumbles with $\tau\to 1$ and does not improve as $(N,T)\to\infty$.
@@
Table 3 below presents the frequency of the correctly selected number of CAs, $g=4$, with the default penalties, $p_{N,T,1}$ and $p_{N,T,2}$ from 2.2, and the penalty from Bai, (2004), $\widetilde{p}_{N,T}$, for the case of stationary factors ($\tau=0$), the mildly non-stationary case ($\tau=0.4$) and the close to local-to-unity process. The idiosyncratic terms $\mathbf{v}_{i,t}$ and $\epsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units. The Supplement presents additional results with weak correlation over time. We start with case of stationary factors established in the literature and the default penalties. In the small $T$ and $N$ case, $T=N=50$, and for stationary factors, we observe a frequency between $16\%$ and $36\%$, a pattern which resembles the results in MW. [^5] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$. If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors. The selection frequency decreases to zero in the case of non-stationary factors. As expected, consistency, illustrated here by a selection frequency of 100%, is quickly achieved when $N$ and $T$ increase. In the non-stationary cases, consistency requires much larger combinations of $(N,T)$. For the extreme case of a process close to local-to-unit, combinations of $(N,T)>300$ are required. Bai, (2004) proposes to adjust the penalty by $\ln(T)$ in the presence of non-stationary factors. The lower panel of Table 3 presents the frequency of the correctly selected number of CAs with a penalty of $\widetilde{p}_{N,T}=\ln(T)\frac{N+T}{NT}\ln(\frac{NT}{N+T})$ for $\mathrm{IC}_{1}$ and $\widetilde{p}_{N,T}=\ln(T)\frac{N+T}{NT}\ln(C^{2}_{N,T})$ for $\mathrm{IC}_{2}$. We observe for all four ICs a selection frequency of 0 for almost all combinations of $N$ and $T$ and $\tau$. Only in the large sample cases of $N=500$ and $T>300$ and only under stationary factors the selection frequency reaches 100%. This strongly suggests that prescriptions from the PC literature do not directly apply to CAs, as ICs are clearly inconsistent, and over-selection should not be heavily penalized.
Figure 1 presents the share of misselected number of cross-sectional averages using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ and $N=T=100$ (”small sample”) and $N=T=500$ (”large sample”), where we increase $\tau$ in $0.01$ steps from 0 to $0.95$. As in Table 3 the idiosyncratics in $\mathbf{x}_{i,t}$ and $\mathbf{v}_{i,t}$ are uncorrelated over time, but weakly correlated across units. For the small sample, the share of misselected number of CAs remains relatively flat and below $20\%$ until $\tau$ reaches levels of around $0.5$, which is the midpoint. Then, it increases to almost $100\%$ with swings showcasing that the criteria become more unstable as $\tau$ increases. As expected, the degree of non-stationarity is much less of a problem for the large sample, and we observe misselection and instability of the criteria only for a high degree of non-stationarity. To shed more light on the importance of the sample size, we present in Figure 2 the share of misselected number of CAs with varying number of time periods from 25 to 500, $\tau=0.9$ and $N=[100,500]$. The share of misselection remains high for $N=100$ for both ICs, but declines for $N=500$. Again, this presents further evidence of the consistency of the ICs, but large samples are required to use them effectively.

### v4

@@ 2 Econometric Setup
$\textit{(b)}\hskip 5.69054pt\textit{Under}\hskip 2.84526ptM\subset M_{0},\hskip 2.84526pt\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\tau b\ln(T)+\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}}^{0})\right)+R_{N,T},$ (2.6)
where $\boldsymbol{\Omega}_{M.M_{0}}^{0}=\operatorname*{plim}_{(N,T)\to\infty}\boldsymbol{\Omega}_{M.M_{0}}$ (positive semi-definite, non-zero), $|R_{N,T}|=o_{p}(1)$ is the remainder, $\lambda_{j}(\mathbf{A})$ is the $j$-th eigenvalue of $\mathbf{A}$, while $b$ is a number of strictly positive eigenvalues.
Proof: Section 3 of the Supplementary material.
In (a), the CAs are able to approximate all $m$ factors under Assumption 4. Then $\overline{\mathbf{Q}}_{M}$ is exactly (log-determinant of) the denominator of the CCE estimator. It uses the fact that under $M_{0}\subset M$ we have $\mathbf{F}=\overline{\mathbf{X}}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}+o_{p}(1)$ for fixed $T$, and by inserting this into $\mathbf{X}_{i}$, we can show that $\ln\mathrm{det}(\overline{\mathbf{Q}}_{M})=\ln\mathrm{det}\left(\boldsymbol{\Sigma}_{\mathbf{v}}\right)+o_{p}(1)$, whose argument is a positive definite matrix under general factors (see also Westerlund,, 2018). Given that $\ln\mathrm{det}(\overline{\mathbf{Q}}_{M_{0}})$ admits the same asymptotic representation, it is natural that $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$ is negligible. Moreover, the rate is already detected by DVS, MW and Bai and Ng, (2002) (for PC) under stationary $\mathbf{F}$, and it remains identical here. This rate is instructive and determines the properties of penalty $p_{N,T}$, which is the key motivation behind the functional forms in (2.2). Specifically, because redundant gmg-m CAs have an asymptotically negligible contribution to the sum of squared residuals, over-specification does not need to be heavily penalized, and so (2.4) becomes
$\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})\right)/p_{N,T}\to_{p}k\cdot(g-m)>0$ (2.7)
as $(N,T)\to\infty$ if $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$. Note that this stands in sharp contrast to findings in Lemma A4 of Bai, (2004), where excess factor estimates contribute non-trivially. Overall, the usual consistency of CAs under general unknown factors prevails when $M_{0}\subset M$ (see also Westerlund,, 2018, or Stauskas,, 2023). However, the result is more nuanced under $M\subset M_{0}$.
In (b), the rank condition in Assumption 4 is not satisfied. This means that CAs are inconsistent for the mgm-g factors. Note that $\overline{\mathbf{Q}}_{M}$ is a quadratic form in $\mathbf{F}$, and we know that $\left\|\mathbf{F}^{\prime}\mathbf{F}\right\|=O_{p}(T^{1+\tau})$ as implied by Assumption 1. However, the objective function offers normalization by $T$ only as the integration order is unknown. Exactly this imbalance causes the divergence of $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}$. A similar situation arises in Lemma A3 of Bai, (2004), where the central step is to detect the sign of divergence, and we follow this route. In (A.9), $b$ is finite ($b<k$), so $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}+\infty$ and therefore
$\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})\right)/\ln(T)\to_{p}\tau b>0,$ (2.8)
@@
as $(N,T)\to\infty$, as required. Two comments are in order. Firstly, neither $\sum_{j:\lambda>0}\ln\left(\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}}^{0})\right)$ nor $R_{N,T}$ are guaranteed to be positive. In addition, in the Supplement, we show that $|R_{N,T}|=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$, which means that for $\tau\approx 1$ and a small $N$, it vanishes slowly as $T\to\infty$. This can induce $\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0$ in small samples. Hence, it may be necessary to rely on large $N,T$ combinations so that $\tau b\ln(T)$ dominates to detect the minimum of IC outside of $M\subset M_{0}$ region if $\mathbf{F}$ is very persistent. We explore and confirm such risks in Monte Carlo simulations in Section 3, where we see that both MW and DVS misselect (in fact, underselect), unless both $N$ and $T$ are large. Secondly, while this result is similar to Lemma A3 in Bai, (2004), there the divergence rate is $O_{p}(T/\ln\ln(T))$, since the law of iterated logarithm is used to arrive at the expression analogous to (A.9) when factors have unit root. Theorem 1 leads to our consistency result.
Theorem 1. Under conditions of Proposition 1 with $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$, we have as $(N,T)\to\infty$
$\mathbb{P}\left(\widehat{M}=M_{0}\right)\to 1.$
@@
Proof. Follows from Proposition 1, because $\mathbb{P}\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})<0\right)=\mathbb{P}\left(\left(\mathrm{IC}(M)-\mathrm{IC}(M_{0})\right)/p_{N,T}<0\right)\to 0$ under (a) and a similar statement holds under (b).
@@
The key message of Theorem 1 is that the IC of DVS (and MW) is consistent, but exhibits hybrid asymptotic properties characteristic to stationary and non-stationary factors. Consequently, two forces are in effect: $p_{N,T}$ should still be negligible, but the factor estimation error can be “large” if $M\subset M_{0}$ as indicated by the remainder in (A.9). To balance these properties in practice, it may be tempting to experiment further with the penalty. Indeed, the change of penalty is the key message of Theorem 1 in Bai, (2004) in the PC setting, where two conditions are satisfied: 1) $p_{N,T}\to\infty$, and 2) $p_{N,T}/\ln(T)\to 0$ which is given in footnote 3 of the study to adapt the penalty to the logarithmic sum of squared residuals, which is exactly our setting. Importantly, this condition does not interfere with consistency in our Theorem 1, because clearly $p_{N,T}=o(\ln(T))$ holds. Therefore, we compare both types of penalties in the simulations to illustrate the outcome if practitioners simply extrapolate recommendations from the PC literature. We let $\widetilde{p}_{N,T}=\ln(T)p_{N,T}$, which follows the suggestion in (12) in Bai, (2004), and obeys $\widetilde{p}_{N,T}\to\infty$ and $\widetilde{p}_{N,T}/\ln(T)\to 0$.
@@
Remark 1. Note that under stationarity, $\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}\to_{p}c>0$ when $M\subset M_{0}$, according to the results in DVS (and similarly in MW). Our Proposition 1 naturally accommodates this result under $\tau=0$, because, by following the same approximation steps, (A.9) becomes
@@
$\mathrm{d}\overline{\mathbf{Q}}_{M,M_{0}}=\sum_{j:\lambda>0}\ln\left(1+\lambda_{j}(\boldsymbol{\Omega}_{M.M_{0}}^{0})\right)+R_{N,T}\to_{p}c>0$
@@
as $(N,T)\to\infty$ because each summand is greater than 1 w.p. 1. Here, $\boldsymbol{\Omega}_{M.M_{0}}^{0}=\operatorname*{plim}_{(N,T)\to\infty}[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}]\overline{\mathbf{Q}}_{M_{0}}^{-1}$ (positive semi-definite, non-zero). Also, $|R_{N,T}|=O_{p}(N^{-1})+O_{p}((NT)^{-1/2})$, which is the same as in DVS (and MW) under $\tau=0$, as expected from the discussion of (A.9).
@@ 3 Simulation Study
where $\mathbf{X}_{i}=\left[\mathbf{x}_{i,1},...,\mathbf{x}_{i,T}\right]^{\prime}$ is a $T\times k$ matrix of observable regressors, $\boldsymbol{\beta}=\iota^{\prime}0.5$ a $k\times 1$ vector of unobserved parameters. The factor loadings are generated as $\boldsymbol{\Gamma}_{i}=[\mathbf{I}_{m},\mathbf{0}_{m\times(k-m)}]\psi_{i}$, that is, a matrix $m\times k$, with $\psi_{i}\sim N(1,1)$ and the elements of the vector $1\times m$ $\boldsymbol{\gamma}_{i}$ are drawn from $N(1,1)$. In the case of errors independent over time and cross-sectionally, $\mathbf{e}_{i,t}$ is drawn from $N(\mathbf{0}_{k+1},\mathbf{I}_{k+1})$. In the case of errors correlated weakly across time and/or cross-sections we follow Bai and Ng, (2002); Margaritella and Westerlund, (2023):
@@
where $\boldsymbol{\epsilon}_{t}$ is a $N\times 1$ stack of $\epsilon_{i,t}\sim N(0,1)$ over $i$, and $\mathbf{w}_{i}$ is the $i$-th row ($1\times N$) of $\mathbf{W}$, a weight matrix with the $J$-th off diagonal elements equal to zero. $\rho$ and $\rho_{v}$ control the correlation over time in $\boldsymbol{\varepsilon}_{t}$ and $\mathbf{V}_{i}$, respectively. For this simulation exercise, we set them to zero. In the Supplement we explore settings with $\rho=\rho_{v}=0.5$, an extension to our theory which prescribes uncorrelated innovations. Across all specifications, we allow for weak correlation across units with $\kappa=\kappa_{v}=0.2$, $J=J_{v}=5$ and $\mathbf{w}_{i}=\tilde{\mathbf{w}}_{i}$. $\mathbf{\Upsilon}_{t}$ is a $N\times k$ matrix, which stacks $\boldsymbol{\upsilon}_{i,t}\sim N(\mathbf{0}_{k},\mathbf{I}_{k})$ over $i$, and it is independent of $\boldsymbol{\epsilon}_{t}$. The vector $\boldsymbol{\iota}_{N,i}\in\mathbb{R}^{N\times 1}$ contains 1 in the $i$-th coordinate, and zeros elsewhere. The factors are drawn as $\mathbf{f}_{t}=\mathbf{R}_{f,T}\mathbf{f}_{t-1}+\mathbf{u}_{f,t}$, where
$\mathbf{R}_{f,T}=\mathbf{I}_{m}-\mathbf{G}T^{-\tau},\quad\mathbf{G}=\mathrm{diag}(g_{1},...,g_{m}),g_{j}\sim U(0,2)\quad\mathbf{u}_{f,t}\sim N\left(\mathbf{0}_{m},\left(\mathbf{I}_{m}-\mathbf{R}_{f,T}^{2}\right)^{1/2}\right).$ (3.3)
@@
In the stationary factors case, the factors are drawn as a stable VAR with $\tau=0$. Two cases of non-stationary factors are considered: the non-stationary case of $\tau=0.4$, and the case close to local-to-unity process $\tau=0.9$. All simulations have 1000 repetitions, and we compare DVS and MW. In the Supplement, we provide additional simulation evidence of the eigenvalue growth ratio (ER) based on Juodis, (2022) for comparison. The focus of this method is the number of factors ($m$), and it does not select the sufficient set of CAs. Nevertheless, its performance crumbles with $\tau\to 1$ and does not improve as $(N,T)\to\infty$.
@@
Table 3 below presents the frequency of the correctly selected number of CAs ($g=4$) with default penalties $p_{N,T,1}$ and $p_{N,T,2}$ from (2.2), and the penalty from Bai, (2004), $\widetilde{p}_{N,T}$, for the case of stationary factors ($\tau=0$), the mildly non-stationary case ($\tau=0.4$) and the close to local-to-unity process. The idiosyncratic terms $\mathbf{v}_{i,t}$ and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated over units. In what follows, the low selection frequency reported is dominated by underslection. We start with the case established in the literature, which is stationary factors and default penalties. For small $N$ and $T$ ($T=N=50$) and stationary factors, we observe a frequency between $16\%$ and $36\%$, a pattern which resembles the results in MW. [^5] In comparison to Margaritella and Westerlund, (2023) we set the number of factors to 4 rather than 2 and hence underestimate the number of CAs more severely in the case of $T=20$ and $N>20$. If we reduce the number of factors to 2, we identify a similar number of factors as in the case of 4 factors. As expected, consistency, illustrated here by a selection frequency of 100%, is quickly achieved when $N$ and $T$ increase. We now turn to the case of non-stationary factors. For small $N$ and $T$, the selection frequency decreases to zero, which is driven by underselection of the number of CAs, as we show in Section 4 in the Supplement. To gain consistency, much larger combinations of $(N,T)$ are required. For the extreme case of a process close to local-to-unity, combinations of $(N,T)>300$ are integral. The behavior of both criteria with respect to increases in $N$, $T$, and $\tau$ is almost identical. Bai, (2004) proposes to adjust the penalty by $\ln(T)$ in the presence of non-stationary factors. The lower panel of Table 3 presents the frequency of the correctly selected number of CAs with a penalty of $\widetilde{p}_{N,T}=\ln(T)\frac{N+T}{NT}\ln(\frac{NT}{N+T})$ for $\mathrm{IC}_{1}$ and $\widetilde{p}_{N,T}=\ln(T)\frac{N+T}{NT}\ln(C^{2}_{N,T})$ for $\mathrm{IC}_{2}$. We observe a selection frequency of 0 for all four ICs and for almost all combinations of $N$ and $T$ and $\tau$. Only in the large sample cases of $N=500$ and $T>300$ and only under stationary factors the selection frequency reaches 100%. This strongly suggests that prescriptions from the PC literature do not directly apply to CAs, as ICs perform even worse, and selecting more CAs should not be heavily penalized.
In the Supplement, we also investigate the share of misselected number of cross-sectional averages when increasing $\tau$ using $IC^{MW}_{1}$ and $IC^{DVS}_{1}$ for a small ($N=T=100$) and large sample ($N=T=500$). For the small sample, the share remains relatively flat until $\tau$ reaches levels of around $0.5$. As expected, this is much less of a problem for the large sample and we observe under-selection of CA only for a high degree of non-stationarity.